# Exploring Mutual Empowerment Between Wireless Networks and RL-based LLMs: A Survey

### v2

@@ Abstract
Reinforcement learning (RL)-based large language models (LLMs), such as ChatGPT, DeepSeek, and Grok-3, have gained significant attention for their exceptional capabilities in natural language processing and multimodal data understanding. Meanwhile, the rapid expansion of information services has driven the growing need for intelligence, efficient, and adaptable wireless networks. Wireless networks require the empowerment of RL-based LLMs while these models also benefit from wireless networks to broaden their application scenarios. Specifically, RL-based LLMs can enhance wireless communication systems through intelligent resource allocation, adaptive network optimization, and real-time decision-making. Conversely, wireless networks provide a vital infrastructure for the efficient training, deployment, and distributed inference of RL-based LLMs, especially in decentralized and edge computing environments. This mutual empowerment highlights the need for a deeper exploration of the interplay between these two domains. We first review recent advancements in wireless communications, highlighting the associated challenges and potential solutions. We then discuss the progress of RL-based LLMs, focusing on key technologies for LLM training, challenges, and potential solutions. Subsequently, we explore the mutual empowerment between these two fields, highlighting key motivations, open challenges, and potential solutions. Finally, we provide insights into future directions, applications, and their societal impact to further explore this intersection, paving the way for next-generation intelligent communication systems. Overall, this survey provides a comprehensive overview of the relationship between RL-based LLMs and wireless networks, offering a vision where these domains empower each other to drive innovations.
Index Terms:
@@
Reinforcement learning (RL), large language models (LLMs), RL-based LLMs, wireless communication.
Figure 1: Illustration of LLM capabilities.
I Introduction
@@
Nowadays, wireless communications [1, 2] have become the backbone of modern society, playing an essential role in a wide range of applications, including edge networks [3, 4, 5], satellite communication [6, 7, 8], the Internet of Things (IoT) [9, 10], drone communication [11, 12], and vehicular networks (V2X) [13, 14]. However, as the demand for efficient information services continues to surge, the pressure on wireless networks has increased, prompting numerous efforts to develop advanced algorithms aimed at alleviating the network burden [15, 16, 17, 18]. Recently, researchers have increasingly leveraged artificial intelligence (AI) to address key challenges in wireless networks, including network performance optimization [19, 20], resource management [21, 22, 23], and the design of efficient semantic communication systems [24, 25]. These advancements underscore AI’s pivotal role in enhancing the efficiency and intelligence of wireless networks.
@@
More recently, significant progress has been made in the field of natural language processing (NLP), with large language models (LLMs) such as ChatGPT [26], Gemini [27], and DeepSeek-R1 [28] gaining substantial attention for their advanced language understanding and human-like text generation capabilities, as shown in Fig. 1. The emergence of these foundational models marks an important milestone in the development of AI, providing highly effective solutions for various complex NLP tasks. For instance, LLMs have been utilized for autonomous code generation [29], medical diagnosis [30], and intelligent tutoring systems [31], demonstrating their versatility across diverse domains. Notably, most state-of-the-art LLMs currently integrate reinforcement learning (RL) [32] to further enhance their performance, leveraging human feedback during the post-training phase [33]. For example, InstructGPT [34], GPT-4 [26], Starling-7B [35], and ChatGLM [36] all employ RL approaches to align their models with human feedback, including human preferences, values, and to reduce hallucinations. Meanwhile, it is worthwhile to note that DeepSeek [28] has made a breakthrough in LLM training by demonstrating that purely RL-based methods, even without any supervised data, can significantly enhance a model’s reasoning capabilities. This advancement highlights the potential of RL in refining LLM performance while reducing dependence on manually labeled datasets.
@@
Beyond NLP, groundbreaking advancements have also emerged in computer vision. Meta AI’s segment anything model (SAM) [37] and its successor, SAM-2 [38], have demonstrated remarkable progress in both image and video processing. These developments underscore the growing convergence of vision and language models, paving the way for more advanced multimodal AI systems. With their strong generalization capabilities and adaptability, these foundational models are being explored for transformative applications across various interdisciplinary fields, including wireless networks. In particular, by leveraging the advantages of RL in long-term optimization, autonomous exploration, and scalability [39, 40, 41], RL-based LLMs present a promising path for optimizing wireless networks, enabling intelligent decision-making, efficient resource allocation, and autonomous adaptation to dynamic environments. For example, [42] adapts LLMs for channel prediction in MISO-OFDM systems, significantly enhancing the accuracy of downlink channel status information prediction through refined training and customized modules. In addition, [43] leverages the RL framework to optimize LLM deployment in edge computing environments, effectively balancing inference performance and computational load by determining the optimal split point.
@@
Despite these promising prospects, integrating RL-based LLMs into wireless networks presents several challenges. First, the dynamic and stochastic nature of wireless environments requires models to effectively balance exploration and exploitation while ensuring real-time adaptability [44, 45]. However, LLMs are well known for being large-scale and resource-intensive, requiring substantial computational resources even during inference. For example, training models such as LLaMA (65 billion parameters) takes 2,048 NVIDIA A100 GPUs running for 21 days [46], while training GPT-3 (175 billion parameters) takes 1,024 NVIDIA V100 GPUs for 34 days [47]. Moreover, inference remains computationally expensive; for instance, when running inference on 128 A100 GPUs, the cost of processing an 8k sequence with GPT-4 is approximately \$0.0049 per 1,000 tokens [48]. Second, optimizing large-scale wireless networks requires efficient resource management, including bandwidth, computational power, energy consumption, and spectrum allocations [49, 50, 45]. However, the computational demands of LLMs are typically associated with high energy consumption, making them challenging to deploy on energy-constrained wireless devices, such as edge devices or IoT nodes. Third, due to the dynamic nature of wireless environments, wireless devices need to adapt to fluctuating network conditions, varying traffic loads, and interference [51, 52]. However, since LLMs are typically pre-trained and difficult to modify, their application in wireless environments requires careful consideration of their ability to adapt to such changes, including changes in channel states and network conditions. Finally, deep learning models are widely known to be vulnerable to various attacks, such as adversarial [53, 54, 55, 56], backdoor [57, 58, 59], and Byzantine attacks [60, 61, 62], which can lead to harmful or incorrect predictions. Therefore, ensuring the robustness and security of RL-LLMs against attacks is crucial for maintaining network reliability.
Figure 2: Organization of the Survey.
@@
On the other hand, wireless networks can also empower RL-based LLMs by serving as a robust infrastructure for efficient training, deployment, and inference, particularly in decentralized and edge computing environments. By leveraging the advantages of wireless networks, RL-based LLMs can achieve real-time data acquisition, low-latency interactions, large-scale distributed training, and scalable model adaptation in dynamic scenarios. For instance, edge computing [63, 9, 64] and federated learning [65, 5, 54, 66] frameworks enable RL-based LLMs to be trained and updated closer to the data source, reducing latency and enhancing privacy by minimizing raw data transmission. This is particularly beneficial in scenarios requiring low-latency decision-making, such as autonomous driving [67], industrial automation [68], and smart healthcare [69, 70]. Moreover, the advancement of high-speed, low-latency wireless technologies can further facilitate efficient model updates and enable large-scale collaborative training across geographically distributed nodes [71].
@@
However, integrating wireless networks into RL-based LLMs also presents several challenges, which can be broadly categorized into four key aspects: data, model, environment, and user. From the data perspective, transmitting privacy-sensitive information, such as user activity or browsing history, over wireless channels increases the risk of data breaches and manipulation [72]. From the model perspective, the LLMs must adapt to fluctuating network conditions, handle limited computational resources at the edge, and mitigate malicious attacks that can manipulate learning processes [73]. Additionally, maintaining model consistency and convergence across decentralized and distributed training settings further complicates deployment [74, 75]. From an environmental perspective, wireless networks often face constraints such as limited bandwidth, computing resources, and energy efficiency [50], which can hinder real-time inference, delay model updates, and introduce trade-offs between performance and resource consumption. Finally, from a user perspective, optimizing RL-based LLMs for diverse user demands and usage scenarios requires adaptive mechanisms that balance personalization, fairness, and efficiency while mitigating potential biases and security threats [76, 77].
@@
In this paper, we provide a comprehensive survey on the mutual empowerment between wireless networks and RL-based LLMs, exploring their motivations, challenges, potential solutions, future directions, and societal impact. To begin, we examine the recent advancements in wireless communication, highlighting the challenges and potential solutions in meeting the growing demands of modern applications. Subsequently, we explore the advancements in RL-based LLMs, offering an in-depth analysis of key training technologies, with a particular focus on the RL strategies used in LLM training, along with the associated challenges and potential solutions. Moreover, we explore the motivations behind the mutual empowerment of wireless networks and RL-based LLMs, illustrating how their synergy can enhance each other’s strengths and drive transformative advancements. We then examine the challenges of integrating advanced LLMs into wireless communication systems and incorporating the benefits of wireless networks into LLM models, followed by insights into effective solutions for both approaches. Furthermore, we discuss potential future directions for deploying RL-based LLMs in wireless communications and introducing wireless communication systems into LLMs, highlighting how emerging technologies can help overcome current limitations. Finally, we explore the broad applications and societal impact of mutual empowerment between these two domains, paving the way for the development of next-generation intelligent communication systems and the evolution of more powerful and adaptable RL-based LLMs. Overall, by examining the challenges of RL-based LLMs and wireless networks and their interconnections, we aim to inspire future research that will drive the advancement of both fields, ultimately leading to a transformative era that pushes the boundaries of state-of-the-art RL-based LLM and wireless network technologies.
I-A Major Contributions
@@
This is the first comprehensive survey that investigates the mutual empowerment of wireless networks and LLMs in terms of motivation, techniques, and applications. Previous surveys have focused on LLM-powered wireless networks from various perspectives, such as resource management [78], network architectures [79], and prompt engineering [80, 81]. However, these prior works have primarily concentrated on the benefits of wireless networks derived from the powerful generalization capabilities of LLMs. In contrast, in this survey, we explore both the mutual empowerment and the symbiotic relationship between wireless networks and LLMs, considering how they can enhance each other’s capabilities and overcome the challenges of each domain. We begin by summarizing the challenges and advancements in both wireless communication and LLMs. Next, we explore the intersection of these two domains, focusing on their motivations and the associated challenges. Subsequently, we offer novel insights and potential solutions to address these challenges. In addition, we examine the applications arising from the intersection of these two domains. Finally, we look ahead to future directions and examine the societal impact of the interplay between these two domains. In summary, the main contributions of this paper are as follows:
@@
    To the best of our knowledge, we are the first to examine the mutual empowerment between RL-based LLMs and wireless networks, providing a comprehensive analysis of their interplay and the way they complement each other.
@@
    First, we review the recent advancements in wireless communications and LLMs independently, with a particular focus on the most state-of-the-art RL-based LLMs.
@@
    We then discuss the motivations, challenges, and potential solutions for integrating RL-based LLMs into wireless networks. On the other hand, the novelty and difference of our paper is the investigation of how wireless networks can be leveraged to enhance the efficiency and performance of RL-based LLMs.
  • $\bullet$
@@
    Finally, we offer a vision for the future directions, applications, and potential societal impacts resulting from the mutual integration between wireless networks and RL-based LLMs.
I-B Organization of This Paper
@@
The remainder of this survey is organized as follows. Section II provides an overview of recent advancements, related technologies, and challenges in wireless networks. Section III examines the latest developments, key technologies, and challenges in RL-based LLMs. Section IV explores how RL-based LLMs can enhance wireless networks, with Section IV-A outlining the motivations, Section IV-B discussing challenges, and Section IV-C presenting potential solutions. Similarly, Section V investigates how wireless networks can empower RL-based LLMs, with Section V-A covering motivations, Section V-B addressing challenges, and Section V-C proposing potential solutions. Additionally, Section VI explores applications arising from the mutual empowerment of wireless networks and RL-based LLMs. Section VII explores future research directions, while Section VIII examines societal impacts. Finally, Section IX concludes the survey. Fig. 2 illustrates the outline of this paper.
Figure 3: Various strategies to improve wireless communication system: Better management of the resource of the current system, improve the transmission ability by improving the current tradition to semantic communication and sensing the environment network to obtain the knowledge related to the physical environment to enhance the beamforming of antennas.
II Overview of Wireless Networks
The amount of data transmitted through wireless communication has increased exponentially in the past couple of years due to various reasons: the increase in the number of connected devices, the requirement for high-quality data, and the need for hyperconnectivity. With the development of advanced applications, such as Metaverse, 3D Online Gaming, Telemedicine, and Autonomous Vehicles, wireless communication is going to face more challenging problems in catching up with user demands due to the enormous data demands in these high-tech applications. Researchers have deployed various approaches from network management for utilizing the currently available resources, semantic communication for eliminating the redundancy information within the data and achieving new state-of-the art communication theory, to integrated sensing and communication for understanding the communication environment and acquiring better communication scenario. In general, various studies have been proposed to improve the serving capacity of the wireless communication ranging from resource management for the existing system, improve the communication design, study the dynamic of the environment to achieve high quality links. Fig. 3 illustrates various strategies for enhancing wireless communication systems.
II-A Network Management
Within the network management approach, a large number of techniques have been considered to facilitate the massive number of connected devices, including advanced modulation and multiple access techniques, edge computing, resource optimization, software-defined networking (SDN), and hybrid network architecture. Each individual technique has its advantages and can complement others in a unique way to achieve an efficient wireless communication framework. For bandwidth resource management, Orthogonal Frequency Division Multiplexing (OFDM)[82, 83, 84] and Non-Orthogonal Multiple Access (NOMA)[85, 86, 87] are two prominent and contrasting methods, OFDM divides the bandwidth spectrum into subcarriers to enable the wireless network to transmit data without interference, while Non-Orthogonal Multiple Access (NOMA) techniques share the same frequency but use power-domain and code-domain for signal separations. Mobile Edge Computing (MEC) [88] was initially proposed to improve the Quality of Service (QoS) for users by offloading computation from centralized cloud servers to edge nodes, reducing latency and enhancing scalability. By deploying computing resources closer to end-users, MEC minimizes communication delays and alleviates network congestion, making it particularly suitable for real-time applications. By decoupling the data and control planes into individuals, the SDN [89] can control the network effectively by adjusting the traffic routing, spectrum allocation to multiple access points, and slicing networks for different applications.
The current wireless communication structure heavily depends on the deployment of fixed base stations, where it encounters difficulties in providing high-quality services for users in rural areas or special occasions such as natural disaster rescue or military operations. To facilitate the need for ubiquitous connectivity, non-terrestrial networks (NTNs) are designed to cover much broader areas[90, 91, 92]. The non-terrestrial architecture can be deployed by using an Unmanned Arial Vehicular (UAV), High Altitude Platform (HAP), or Satellites (Low-earth Orbit, Medium-earth Orbit, Geostationary Orbit). The non-terrestrial network can be established by deploying one type of device mentioned above or combinations of different types of devices, each of which offers unique benefits; the UAVs offer flexibility for the network based on ease of flight path adjustment, while the satellites extend the coverage area to much bigger regions and more resources (computing, energy) compared to UAV. With the moving property, the NTNs can create communication connections to remote users that the fixed base station (BS) cannot reach. However, the proposal of NTNs, along with the MEC, increases the complexity of the wireless communication infrastructure, and it needs to be managed efficiently to be able to exploit the proposal’s capabilities fully [93, 94, 95].
II-B Semantic Communication

### v3

# DeepSeek-Inspired Exploration of RL-based LLMs and Synergy with Wireless Networks: A Survey

@@ Abstract.
Recently, reinforcement learning (RL)-based large language models (LLMs), such as ChatGPT, DeepSeek, and Grok-3, have attracted significant attention for their impressive capabilities in multimodal data understanding. Meanwhile, as information services rapidly expand, there is increasing demand for intelligent and adaptable wireless networks to support this growth. Integrating LLMs with wireless networks offers a promising solution, as LLMs enhance network optimization with strong reasoning and decision-making abilities, while wireless infrastructure enables broad deployment of these models. This synergy calls for deeper exploration into the convergence of the two fields. To this end, this paper makes a timely contribution by first reviewing key technologies for wireless network optimization, establishing a foundation for understanding how LLMs can be effectively integrated. We then turn to recent advancements in RL-based LLMs, particularly the open-source DeepSeek models, as they offer greater accessibility and customizability for researchers and practitioners. Subsequently, we explore the synergy between these fields, emphasizing motivations, open challenges, and potential solutions. Finally, we provide insights into future directions, applications, and societal impacts. Overall, this survey offers a comprehensive exploration of the relationship between LLMs and wireless networks, presenting a vision of how these domains empower each other to drive innovation.
@@
Reinforcement learning, large language models, RL-based LLMs, wireless networks
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Computing methodologiesccs: Natural language generationccs: Artificial intelligenceccs: Machine learning
1. Introduction
@@
Wireless communications (Liu et al., 2024a; Nguyen et al., 2024c) have become the backbone of modern society, playing an essential role in a wide range of applications, including edge networks (Qiao et al., 2025a; Xu et al., 2023; Qiao et al., 2024a), satellite communication (Talgat et al., 2024; Fourati and Alouini, 2021; Hassan et al., 2024a), Internet of Things (IoT) (Qiao et al., 2023b; Zhang et al., 2015), drone communication (Mozaffari et al., 2018; Hayajneh et al., 2016), and vehicular networks (V2X) (Gerla and Kleinrock, 2011; Nguyen et al., 2024b). However, as the demand for efficient information services continues to surge, the pressure on wireless networks has increased, prompting numerous efforts to develop advanced algorithms aimed at alleviating the network burden (Jin and Qiao, 2020; Cichoń et al., 2016; Ge et al., 2024; Zou et al., 2024). Recently, researchers have increasingly leveraged artificial intelligence (AI) technologies to address key challenges in wireless networks, including efficient resource management (Adhikary et al., 2023a; Khoramnejad and Hossain, 2025; Raha et al., 2023b), intelligent network architecture (Adhikary et al., 2024c; Cao et al., 2017b; Adhikary et al., 2023b), and the design of seamless connectivity and quality service (Raha et al., 2023a; Nguyen et al., 2025a). These advancements underscore AI’s pivotal role in enhancing the efficiency and intelligence of wireless networks.
@@
More recently, significant progress has been made in the field of natural language processing (NLP), with large language models (LLMs) such as ChatGPT (Achiam et al., 2023), Gemini (Team et al., 2023), and DeepSeek (Guo et al., 2025) gaining substantial attention for their advanced language understanding and human-like text generation capabilities. The emergence of these foundational models marks an important milestone in the development of AI, providing highly effective solutions for various complex NLP tasks. For instance, LLMs have been utilized for autonomous code generation (Dong et al., 2024), medical diagnosis (Garg et al., 2023), and intelligent tutoring systems (Vujinović et al., 2024), demonstrating their versatility across diverse domains. Notably, most state-of-the-art LLMs currently integrate reinforcement learning (RL) (Milani et al., 2024) to further enhance their performance by leveraging human feedback during the post-training phase (Wang et al., 2024e). Notable examples include InstructGPT (Ouyang et al., 2022), GPT-4 (Achiam et al., 2023), Starling-7B (Zhu et al., 2024a), and ChatGLM (GLM et al., 2024), all of which incorporate RL to align outputs with user preferences better and reduce hallucinations. Among them, DeepSeek stands out as a pioneering effort demonstrating that purely RL-based training, even without supervised data, can significantly improve LLM reasoning capabilities (Guo et al., 2025). Moreover, its open-source release lowers the barrier to entry for researchers and practitioners, offering greater accessibility, transparency, and customizability. Motivated by these unique features, this paper adopts DeepSeek as a representative and inspirational case study to explore the broader landscape of RL-based LLMs.
@@
Beyond NLP, groundbreaking advancements have also emerged in computer vision. Meta AI’s segment anything model (SAM) (Kirillov et al., 2023) and its successor, SAM-2 (Ravi et al., 2024), have demonstrated remarkable progress in both image and video processing. These developments underscore the growing convergence of vision and language models, paving the way for more advanced multimodal AI systems. With their strong generalization capabilities and adaptability, these foundational models are being explored for transformative applications across various interdisciplinary fields. In the context of wireless networks, leveraging the advantages of RL in long-term optimization, autonomous exploration, and scalability (Lei et al., 2020; Ladosz et al., 2022; Ei et al., 2024), RL-based LLMs offer a promising path for optimizing wireless networks. They can enable intelligent decision-making, efficient resource allocation, and autonomous adaptation to dynamic environments (Chen et al., 2024c). For instance, (Sevim et al., 2024) incorporates RL into the training of an LLM-assisted agent, enabling it to determine the optimal placement and orientation of a base station in urban environments to maximize coverage. In addition, (Chen et al., 2024c) leverages the RL framework to optimize LLM deployment in edge computing environments, effectively balancing inference performance and computational load by determining the optimal split point.
@@
Despite these promising prospects, integrating RL-LLMs into wireless networks still presents several challenges. First, the dynamic and stochastic nature of wireless environments requires models to effectively balance exploration and exploitation while ensuring real-time adaptability (Bai et al., 2023; Wang et al., 2024b). However, LLMs are well known for being large-scale and resource-intensive, requiring substantial computational resources even during inference. For example, training models such as LLaMA (65 billion parameters) takes 2,048 NVIDIA A100 GPUs running for 21 days (Touvron et al., 2023), while training GPT-3 (175 billion parameters) takes 1,024 NVIDIA V100 GPUs for 34 days (Zhao, 2024). Moreover, inference remains computationally expensive; for instance, when running inference on 128 A100 GPUs, the cost of processing an 8k sequence with GPT-4 is approximately \$0.0049 per 1,000 tokens (Wallstreetcn, 2023). Second, optimizing large-scale wireless networks requires efficient resource management, including bandwidth, computational power, energy consumption, and spectrum allocations (Chen et al., 2024a; Cao et al., 2017a; Wang et al., 2024b). However, the computational demands of LLMs are typically associated with high energy consumption, making them challenging to deploy on energy-constrained wireless devices, such as edge devices or IoT nodes. Third, due to the dynamic nature of wireless environments, wireless devices need to adapt to fluctuating network conditions, varying traffic loads, and interference (Letaief and Zhang, 2006; Ji et al., 2015). However, since LLMs are typically pre-trained and difficult to modify, their application in wireless environments requires careful consideration of their ability to adapt to such changes, including changes in channel states and network conditions. Finally, deep learning models are widely known to be vulnerable to various attacks, such as adversarial (Qiao et al., 2024c; Goodfellow et al., 2015), backdoor (Liu et al., 2020; Gao et al., 2020; Doan et al., 2021), and Byzantine attacks (Cui et al., 2024; Li et al., 2024a; Wang et al., 2024c), which can mislead the learning process or manipulate model outputs. In RL-based LLMs, the iterative interactions with the environment or humans make them particularly vulnerable to harmful feedback or adversaries. Therefore, ensuring the robustness and security of RL-LLMs against attacks is also a crucial concern for maintaining network reliability.
Figure 1. Organization of the Survey.
@@
On the other hand, wireless networks can also empower LLMs by serving as a robust infrastructure for efficient training, deployment, and inference, particularly in decentralized and edge computing environments. By leveraging the advantages of wireless networks, LLMs can achieve real-time data acquisition, low-latency interactions, large-scale distributed training, and scalable model adaptation in dynamic scenarios. In particular, RL-based strategies for training LLMs are well-suited to such environments, as their effectiveness relies on continuous interaction with the environment and the ability to adapt in response to real-time feedback. For instance, edge computing (Shi et al., 2016; Qiao et al., 2023b, d) and federated learning (McMahan et al., 2017; Qiao et al., 2024a, d; le G et al., 2024) frameworks can bring RL-LLMs closer to the data source, enabling timely updates, reduced communication latency, and enhanced privacy by minimizing raw data transmission. This is particularly beneficial in scenarios requiring low-latency decision-making, such as autonomous driving (Yang et al., 2021), industrial automation (Pandey et al., 2024), and smart healthcare (Lai and Lai, 2022; El Zouka and Hosni, 2021). Moreover, the advancement of high-speed, low-latency wireless technologies can further facilitate efficient model updates and enable large-scale collaborative training across geographically distributed nodes (Javaid et al., 2024).
@@
However, integrating wireless networks into RL-based LLMs also presents several challenges, which can be broadly categorized into four key aspects: data, model, environment, and user. From the data perspective, transmitting privacy-sensitive information, such as user activity or browsing history, over wireless channels increases the risk of data breaches and manipulation (Guo et al., 2024). From the model perspective, the LLMs must adapt to fluctuating network conditions, handle limited computational resources at the edge, and mitigate malicious attacks that can manipulate learning processes (Qiao et al., 2024b). Additionally, maintaining model consistency and convergence across decentralized and distributed training settings further complicates deployment (Qiao et al., 2023c; Shi et al., 2023). From an environmental perspective, wireless networks often face constraints such as limited bandwidth, computing resources, and energy efficiency (Cao et al., 2017a), which can hinder real-time inference, delay model updates, and introduce trade-offs between performance and resource consumption. Finally, from a user perspective, optimizing LLMs for diverse user demands and usage scenarios requires adaptive mechanisms that balance personalization, fairness, and efficiency while mitigating potential biases and security threats (Zhao et al., 2025b; Zhu et al., 2024c).
@@
Building on the previous discussion, this paper presents a comprehensive survey on the synergy between wireless networks and RL-based LLMs, exploring their motivations, challenges, potential solutions, future directions, and societal impact. To begin, we examine the recent advancements in wireless networks, highlighting the challenges and potential solutions in meeting the growing demands of modern applications. Subsequently, we explore the advancements in LLMs, offering an in-depth analysis of key training technologies, with a particular focus on the RL strategies used in LLM training, along with the associated challenges and potential solutions. Moreover, we explore the motivations behind the mutual empowerment of wireless networks and RL-based LLMs, illustrating how their synergy can enhance each other’s strengths and drive transformative advancements. We then examine the challenges of integrating advanced LLMs into wireless communication systems and incorporating the benefits of wireless networks into LLM models, followed by insights into effective solutions for both strategies. Furthermore, we discuss potential future directions for deploying LLMs in wireless network environments and introducing wireless communication systems into LLMs, highlighting how emerging technologies can help overcome current limitations. Finally, we explore the broad applications and societal impact of the synergy between these two domains, paving the way for the development of next-generation intelligent communication systems and the evolution of more powerful and adaptable LLMs. Overall, by examining the challenges of LLMs and wireless networks and their interconnections, we aim to inspire future research that will drive the advancement of both fields, ultimately leading to a transformative era that pushes the boundaries of state-of-the-art LLMs and wireless network technologies.
1.1. Major Contributions
@@
This is the first comprehensive survey that investigates the synergy between wireless networks and RL-based LLMs in terms of motivation, techniques, and applications. Previous surveys have focused on LLM-powered wireless networks from various perspectives, such as resource management (Boateng et al., 2024; Adhikary et al., 2024c), network architectures (Friha et al., 2024; Xu et al., 2024b), and prompt engineering (Zhou et al., 2024a; Shao et al., 2024b). However, these prior works have primarily concentrated on the benefits of wireless networks derived from the powerful generalization capabilities of LLMs. In contrast, in this survey, we explore both the mutual empowerment and the symbiotic relationship between wireless networks and LLMs, considering how they can enhance each other’s capabilities and overcome the challenges of each domain. We begin by summarizing the challenges and advancements in both wireless communication and LLMs. Next, we explore the intersection of these two domains, focusing on their motivations and the associated challenges. Subsequently, we offer novel insights and potential solutions to address these challenges. In addition, we examine the applications arising from the intersection of these two domains. Finally, we look ahead to future directions and examine the societal impact of the interplay between these two domains. In summary, the main contributions of this paper are as follows:
@@
    Recognizing that prior work has mainly focused on the role of LLMs in enhancing wireless networks, this survey shifts the focus towards how wireless networks can, in turn, support LLMs. Meanwhile, we revisit the role of LLMs in optimizing wireless networks to provide a more comprehensive understanding.
@@
    Building on a comprehensive review of the key enabling technologies in both domains, we further analyze their synergy by identifying core motivations, major challenges, and potential solution strategies from two complementary perspectives: wireless networks for LLMs and LLMs for wireless networks.
@@
    Grounded in this dual-perspective analysis, we discuss future directions, emerging applications, and potential societal impacts resulting from the mutual empowerment between wireless networks and LLMs.
1.2. Organization of This Paper
@@
The remainder of this survey is organized as follows. Section 2 provides a brief overview of the intersection between wireless networks and LLMs. Specifically, Sections 2.1 and 2.2 introduce the key technologies behind each domain, while Section 2.3 briefly introduces their mutual benefits. Section 3 explores how LLMs can enhance wireless networks, with Section 3.1 outlining the motivations, Section 3.2 discussing challenges, and Section 3.3 presenting potential solutions. Similarly, Section 4 investigates how wireless networks can empower LLMs, with Section 4.1 covering motivations, Section 4.2 addressing challenges, and Section 4.3 proposing potential solutions. In addition, Section 5 explores applications arising from the mutual empowerment of wireless networks and RL-based LLMs. Section 6 explores future research directions, while Section 7 examines societal impacts. Finally, Section 8 concludes the survey. Fig. 1 illustrates the outline of this paper.
Figure 2. Various strategies for enhancing wireless communication systems: optimizing resource management in existing networks, improving transmission efficiency by transitioning from traditional to semantic communication, and leveraging environmental sensing to acquire physical context information for enhanced antenna beamforming.
2. Bridging Wireless Networks and Large Language Models
2.1. Key Technologies for Optimizing Wireless Networks
The volume of data transmitted through wireless communication has surged exponentially in recent years, driven by several factors: the growing number of connected devices, the demand for high-quality data, and the increasing need for hyperconnectivity (Nguyen et al., 2024c; Mozaffari et al., 2018). With the emergence of advanced applications such as the Metaverse (Sami et al., 2024), 3D online gaming (Wiboolyasarin et al., 2024), telemedicine (Lu et al., 2024b), and autonomous vehicles (Parekh et al., 2022), wireless communication systems have been facing increasing challenges in meeting user demands due to the massive data requirements of these high-tech applications (Aceto et al., 2019). To address these challenges, advancements in wireless technologies can be broadly categorized into three complementary areas: (1) efficient resource management, (2) intelligent network architecture, and (3) integration of computing and communication. The resource management approach focuses on optimizing existing infrastructure to accommodate the growing number of devices, employing techniques such as dynamic spectrum allocation (Tun et al., 2019). This method maximizes bandwidth efficiency by allocating more resources to active devices while reducing or suspending allocation for idle ones (Tun et al., 2020). The intelligent network architecture approach explores new transmission paradigms, AI-driven optimization, and security enhancements to improve network performance (Lu et al., 2024a). Lastly, the integration of computing and communications approach investigates cutting-edge physical technologies, including higher-frequency communication, advanced antenna systems, and emerging innovations such as reconfigurable intelligent surfaces (RIS) (Kim et al., 2023). A visual overview of these three solutions is presented in Fig. 2.
2.1.1. Efficient Resource Management
Within the network management approach, various techniques have been explored to accommodate the growing number of connected devices (Cho et al., 2010; Zhang, 2019; Afolabi et al., 2018). These include advanced modulation and multiple access techniques (Dai et al., 2015; Saito et al., 2013), resource optimization strategies (Niu et al., 2024b; Kim et al., 2023), software-defined networking (SDN) (Afolabi et al., 2018), and hybrid network architectures (Fu et al., 2024). Each individual technique has its advantages and can complement others in a unique way to achieve an efficient wireless communication framework.
Communication Resource Management. Orthogonal Frequency Division Multiplexing (OFDM) (Nee and Prasad, 2000; Cho et al., 2010; Stuber et al., 2004) and non-orthogonal multiple access (NOMA) (Saito et al., 2013; Dai et al., 2015; Kim and Lee, 2015) represent two prominent yet contrasting approaches. OFDM partitions the bandwidth spectrum into subcarriers, allowing wireless networks to transmit data without interference. In contrast, NOMA enables multiple users to share the same frequency by leveraging power-domain and code-domain techniques for signal separation. Another advancement in this approach is network slicing, which involves partitioning the network into several logical segments to support distinct application scenarios, such as enhanced mobile broadband (eMBB) (Tun et al., 2020), ultra-reliable low-latency communications (URLLC) (Tun et al., 2020), and massive machine-type communications (mMTC) (Dutkiewicz et al., 2017), rather than relying on a one-size-fits-all solution (Zhang, 2019). Reserving bandwidth exclusively for certain applications such as television broadcasting, satellite communication(Sat, [n. d.]), and aeronautical can lead to inefficient use of wireless communication resources. To address this issue, dynamic spectrum access (DSA) has been introduced as a way to make use of underutilized spectrum without interfering with primary users (Zhao and Sadler, 2007). It works by actively identifying idle licensed channels and temporarily assigning them to secondary users. With the help of smart allocation algorithms, DSA offers a promising solution to the growing challenge of spectrum scarcity (Tan et al., 2022).
Network Slicing. Network slicing enables the network to adapt to users’ changing requirements (Lei et al., 2022) dynamically. Note that the underlying concept of network slicing is virtualization, which decouples the network from its hardware dependencies and leverages available resources to run software functions (Afolabi et al., 2018). In this context, virtual machines provide a platform for guest operating systems, enabling applications and services to be executed effectively (Tun et al., 2019, 2020).
2.1.2. Towards Intelligent Network Architecture Design: The Role of Semantic Communication and ISAC
To meet the demands of increasingly complex applications and the explosive growth of connected devices, future wireless networks must evolve beyond rigid, manually managed infrastructures. Intelligent and programmable architectures offer enhanced adaptability, efficiency, and context-awareness, enabling the network to respond dynamically to complex and changing environments. Achieving this vision requires a fundamental redesign of existing network architectures, moving away from traditional network architecture towards deeply integrated and AI-empowered frameworks (Yang et al., 2023a; Xu et al., 2024a; Zhao et al., 2024b). Among them, semantic communication and integrated sensing and communication (ISAC) have emerged as representative innovations that significantly reshape the communication paradigm at both architectural and functional levels.